OpenStudy (anonymous):

Bit confused.... Why do I divide by 3? 3|x + 6| = 36

OpenStudy (anonymous):

You would multiply 3 to x and 6 3x+18=36 minus 18 3x=18 divide by 3 x=6

OpenStudy (swissgirl):

disagree First divide by 3

OpenStudy (anonymous):

wait i dont think thats how you do it

OpenStudy (anonymous):

why do you divide?

OpenStudy (swissgirl):

ok soooo we are trying as much as possible to alienate the absolute value

OpenStudy (swissgirl):

and then once we have just the absolute value isolated on one side of the equation we can the solve for absolute value regularly such as (x+6)=13 and (x+6)=-13

OpenStudy (anonymous):

Dividing simplifies the equation and it's calculations. P.S. it's 12 not 13.

OpenStudy (anonymous):

You apply the opposite operations when you move the numbers/variables to the other side. So, 3x+6=36 3x= 36-6 3x=30 x=30/2 x=15. Get it?

OpenStudy (anonymous):

why isn't it 36 minus 3? instead of division?

OpenStudy (swissgirl):

Sorry you are correct: (x+6)=12 and (x+6)=-12

OpenStudy (anonymous):

lots of different answers.....

OpenStudy (anonymous):

so whatever number is to the left of an absolute value you ALWAYS divide to the other side of the equal sign?

OpenStudy (anonymous):

\[\left| x+6 \right|=\pm \left( x+6 \right)\] solve taking both cases.

OpenStudy (anonymous):

but why do you divideeeeeeeeeeeeeee

OpenStudy (amistre64):

you divide because 3/3 = 1

OpenStudy (anonymous):

but i thought you can only add or subtract the number from the left of the equal sign to the right of the equal sign

OpenStudy (amistre64):

\[3a = b\] \[\frac{3}{3}a = \frac{b}{3}\] \[1a = \frac{b}{3}\] \[a = \frac{b}{3}\]

OpenStudy (amistre64):

no, you can work any mathimatical operation you want to.

OpenStudy (amistre64):

the idea is to use the same operation on both sides in order to keep things equal.

OpenStudy (anonymous):

can you multiply too?

OpenStudy (amistre64):

is multiplication a mathimatical operation?

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

then yes ....

OpenStudy (amistre64):

dividing by 3 is the same as multiplying by 1/3

OpenStudy (anonymous):

*sigh* im so confused right now i thought in algebra you cant divide the other side unless it was like a 3x

OpenStudy (amistre64):

it is 'like' 3x

OpenStudy (amistre64):

3|x+6| realise that |x+6| is just some number ... let it be equal to a 3a

OpenStudy (anonymous):

hmmmm so okay this problem wont work without division?

OpenStudy (amistre64):

itll work with any mathimatical operation you want to perform. but dividing off the 3 is the fastest way to approach it

OpenStudy (amistre64):

we could multiply by 2 if we wanted to, but it doesnt really do much to help us solve the problem in a timely fashion

OpenStudy (amistre64):

we could add 5/16 but that doesnt really do us much in the way of getting to a solution quickly

OpenStudy (anonymous):

its 3 because its the lowest common denominator right?

OpenStudy (amistre64):

its 3 because you have: 3, times, something if we had 1, times, something ... things would move quicker for us since 3/3 = 1, we should divide it all by 3

OpenStudy (anonymous):

what if it was 25|x + 6| = 30

OpenStudy (anonymous):

would it be 5|x + 6| = 6

OpenStudy (amistre64):

thats fine too, or 25, times, something this would be easier if we had: 1, times, something since 25/25 = 1, we should do what?

OpenStudy (amistre64):

30/25 = 6/5 so either approach you take

OpenStudy (anonymous):

doesnt it have to be 5 on each side?

OpenStudy (anonymous):

since thats the lowest multiple?

OpenStudy (amistre64):

no, it doesnt 'have to be'

OpenStudy (amistre64):

nowhere in this conversation has anyone mentioned that you have to use some lowest multiple.

OpenStudy (amistre64):

all youve done is modify the form to: 5|x+6| = 6 you still have to divide off another 5

OpenStudy (anonymous):

so you have to carry the 25 and divide it to 30?

OpenStudy (anonymous):

and you divide 25 by 25?

OpenStudy (anonymous):

so it would be |x +6| = 30/25?

OpenStudy (amistre64):

you want to create the situation where you have: 1|x+6| on the left side

OpenStudy (amistre64):

so dividing 25 by 25 gets us to that situation.

OpenStudy (anonymous):

ohhhhhhhhhhh okay i get it

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

good :) youre welcome